Find each of the function below, indicate whether the function in onto, on-to-one neither or both. If the function is not onto or nor one-to-one, give an example showing why

                               .           

H;Z     Z. h(x)=x^3


1
Expert's answer
2021-07-07T15:06:44-0400

Notice that this function passes BOTH a vertical line test and a horizontal line test.


"y_1=y_2=>x_1^3=x_2^3"

"=>(x_1-x_2)(x_1^2+x_1x_2+x_2^2)=0"

"=>x_1=x_2, \text{where} \,x_1,x_2\in \R"

The function "g(x)=x^3" is one-to-one.


"\forall y\in \R \ \exist x=\sqrt[3]{y}" ​such that "y=x^3."

The function "g(x)=x^3" is onto.


Therefore the function "g(x)=x^3" is bijection.